# Spike-frequency and h-current based adaptation are dynamically equivalent in a Wilson-Cowan field model

### v2

@@ Abstract
Keywords: Wilson-Cowan, neural field model, spike-frequency adaptation, h-currents, symmetry, equivalence transformation, adaptation-driven slow waves
@@ I Introduction
One frequently used model is the Wilson-Cowan model which was first introduced in [33] in its mass-model version, i.e. without a spatial domain. The mass model version displays a rich variety of emergent dynamics and has previously been used to investigate SFA-driven SOs, either with [34, 35] or without inhibition [36]. In its original form, comprising of one excitatory and one inhibitory population of neurons without SFA [37], the dynamical landscape of this model remains fairly similar when both SFA and inhibition are included, when only SFA is present, or when both are absent. A regime of oscillations (e.g., emerging due to a Hopf instability) is at the corner end of a regime of bistability, where a stable up- and down-state coexist.
@@
To include a spatial domain, Wilson and Cowan extended the mass model to a field model which is comprised of two coupled populations of excitatory and inhibitory neurons [38]. Pinto and Ermentrout [39] investigated the system without an inhibitory population, but with a linear, negative feedback that limits the networks excitability. They establish the existence of traveling pulses in such a network and connect the speed of the waves, which are induced by the slow negative feedback, to the model parameters using the Heaviside function as transfer function. A sigmoidal transfer function speeds up waves further when compared to the Heaviside transfer function. For a homogeneous, excitatory neural field model, [40] proved that lateral inhibition is required to stabilize bump solutions (i.e., stabilize the activity to not decay or vanish over time), and while [41] confirms these findings, they additionally show that local inhomogeneous input stabilizes stationary-pulse solutions in an otherwise homogeneous network. In [42], the authors investigate the same field model without inhibition in which nonlinear negative feedback to an excitatory population of neurons is produced by synaptic depression and SFA simultaneously. They identify traveling fronts and pulses, as well as stationary or homogeneous oscillations. They show that synaptic depression is not sufficient to stabilize bump solutions and confirm that either lateral inhibition (as shown in [40]) or external input (as shown in [41]) is required. Conversly, bumps do not exist when SFA is present. Furthermore, they investigated the phase space with respect to the strengths of SFA and synaptic depression and found that SFA has little to no effect on the speed and width of traveling fronts and pulses. Their nonlinear formulation of SFA feedback is a physiologically based negative feedback. In [43], the authors equip the excitatory population of the two-population model with SFA and investigate the conditions for Turing- and Hopf bifurcations to emerge, enabling stationary or spatio-temporal activity. While weaker and faster adaptation generates stationary patterns, slower and stronger adaptation causes the stationary patterns to travel.
@@
Now, we are left with the question of what the differences are between the slow negative feedback mechanism based on SFA, see [44, 45, 46, 47, 35, 48], and the recently proposed positive feedback mechanism based on h-currents, see [21, 22] which have already been shown to induce SO-like dynamics. Simultaneous investigations of h-currents and SFA were already conducted. [49] used fluorescent dye in the dentate gyrus of rats and unmodified versus blocked HCN-channel recordings to show that in the presence of HCN channels action potential (AP) initiation is enhanced during sustained firing and the propagation of AP patterns is sped up. They observe that in the presence of HCN channels, hyperpolarizing Na+-K+ currents are opposed. [50] included both HCN (h-current based adaptation) and M-channel (SFA) dynamics in a spiking neural network to investigate their role in seizure events in epilepsy patients. [22] found, in slice experiments, that blocking h-currents results in prolonged up- and down-state durations, effectively decreasing the temporal oscillation frequency. With their computational model they produce adaptation-driven SOs between up- and down-states where the implemented h-current and SFA based adaptation alternate in activation, depending on the membrane potential with h-currents active in down-, and inactive in up-states while the opposite is the case for SFA. With the model, they recreated the results found in vitro, prolonging up- and down-state durations, if the strength of h-currents is further reduced. [21] tested whether a spatial gradient in excitability, degree of recurrent excitation, or strength of the SFA current would induce a sequential activation of cortical neurons (i.e, up-state onset) along the dorso-ventral axis of the rat post-subiculum with a synchronous deactivation of the up-state. They observed this in vivo in naturally sleeping mice prior to their in-silico experiments. Neither gradient was able to replicate the experimentally observed dynamics. Including h-currents, on the other hand, with decreasing strength along the dorso-ventral axis replicated their results. A detailed analysis of the two opposing mechanisms shaping the spatio-temporal neural dynamics, however, has not yet been performed.
@@ II The spatially extended adaptive Wilson-Cowan model
We use the original Wilson-Cowan continuum model of [38] with integro-differential equations, including a convolution over a spatial domain, to investigate adaptation-driven traveling waves of SO-like dynamics. We include a third differential equation for an adaptation mechanism and connect it to the excitatory population to describe the effects of either high-activity induced negative or low-activity induced positive feedback mechanisms, respectively. In the following we refer to them as spike-frequency ($a$) or h-current ($h$) mediated adaptation although the model is a simplification that focuses only on their most prominent biophysical features. The equations of the activities $u_{e}$ and $u_{i}$ of the excitatory and inhibitory population and for the adaptation variable $m$ are given by
@@
The coupling strengths between the excitatory and inhibitory populations are given by connectivity weights $w_{jl},\ j,l\in\{e,i\}$, in Equation (1). The membrane time constants are designated with $\tau_{j},\ j\in\{e,i\}$. The adaptation mechanisms $m(x,t)$ are discussed in detail in Section II.1; they are parametrized by a strength value $b$, a time constant $\tau_{m}$, and a gain parameter $\beta_{m}$. Due to the transfer function in Equation (4) being a sigmoid between zero and one, we interpret the activity variables $u_{e}$ and $u_{i}$ as the proportion of neurons firing per unit time [54]. Its gain parameter is denoted by $\beta_{j}$. The kernel widths in Equation (3) determine the extent to which populations are effectively connected in space. Kernel widths are chosen to be smaller for excitatory and broader for inhibitory interactions, effectively implementing a Mexican-hat shaped interaction between different locations. The parameters that are used in this study were adjusted such that adaptation-driven SOs are generated with an additional noisy external input (see [46, 34]), fast excitation-inhibition driven oscillations can emerge (see [55]), and the system with and without adaptation shows regimes of bistability since this has been shown to play a crucial part in the formation of SOs (see [51, 56]).
Table 1: List of model parameters.
@@
$w_{ee}$ 3.2 Coupling strength excitatory to excitatory
$w_{ei}$ 2.6 Coupling strength inhibitory to excitatory
$w_{ie}$ 3.3 Coupling strength excitatory to inhibitory
$w_{ii}$ 0.9 Coupling strength inhibitory to inhibitory
$I_{e}$ [-2.5, 2] External input current to exc population
$I_{i}$ [-2.5, 2] External input current to inh population
$\tau_{e}$ $10$ Membrane time constant of exc population
$\tau_{i}$ $15$ Membrane time constant of inh population
$\sigma_{e}$ 1 Spatial spread of exc population
$\sigma_{i}$ 3 Spatial spread of inh population
$\mu$ $0.3$ Adaptation threshold
@@
$b$ $[0,1]$ Spike-frequency adaptation strength
$\tau_{m}$ $300$ Spike-frequency adaptation time constant
$\beta_{m}$ $10$ Spike-frequency adaptation gain
@@
The model is implemented as a 1-dimensional neural field with periodic boundary conditions and a homogeneous connectivity profile. To reduce the effects of the finite size on the dynamics, we combine a large ring of length $L=600$ with comparably small widths $\sigma_{l},\ l\in\{e,i\}$, of the spatial kernels. The ring is discretized into $n$ positions, each of which is equipped with a node whose dynamics is described by Equations (1). Therefore, the integration space step is given by $dx=\frac{L}{n}$.
@@ II.1 The adaptation mechanisms
The excitatory populations are equipped with an adaptation mechanism, generally understood as a process (not restricted to the brain) that modulates the effect of a sustained input but acts on a slower time scale (see [57]). In the following, we distinguish two adaptation processes caused by different types of somatic conductances in cortical neurons (see [58]). Spike-frequency adaptation (SFA), a negative feedback mechanism, is mediated by slow potassium currents which activate when the activity of the excitatory population is high. This causes a hyperpolarizing feedback current which effectively inhibits the neural response of excitatory neurons (see [47]). H-current based adaptation is mediated by hyperpolarization-activated nucleotide-gated channels (see [59]) which open if the excitatory population is in a low-activity state and return positive feedback, i.e., effectively excite the excitatory population. Functionally, the adaptation mechanisms are similar (see [60] for SFA, for h-currents see [61] where the authors discuss h-currents in thalamic neurons, and see [21, 58] for both adaptation mechanisms in cortical neurons), so that we can describe their main effects by the same differential equation (last equation in Equations (1), see [21]) and include the inverse mode of action (positive versus negative feedback, activated during high- versus low-activity values) by changing the sign of the parameters for adaptation strength $b$ and adaptation gain $\beta_{m}$. Parameters are summarized in Table 1.
@@ II.2 Fixed points and stability analysis of the model with adaptation
To determine the fixed points we consider a single Wilson-Cowan node. Due to the normalization of the kernel, its fixed points are equivalent to the fixed points of the spatially extended model. Let $\frac{\partial u_{e}}{\partial t}=\frac{\partial u_{i}}{\partial t}=0$ in Equations (1). From $\frac{\partial m}{\partial t}=0$ we obtain $m=F_{j}(u_{e})$ for $j\in\{a,h\}$ which we insert in the equation for the excitatory population. The nullclines $f(u_{e})$ and $g(u_{i})$ are then given by
$u_{i}=f(u_{e})=$ $\frac{1}{w_{ei}}\bigl(w_{ee}u_{e}+I_{e}-bF_{j}(u_{e})-F^{-1}_{e}(u_{e})\ \bigr),\ j\in\{a,h\},$
@@
A point $(\tilde{u}_{e},\tilde{u}_{i},\tilde{m})\in(0,1)^{3}$ is a fixed point if $f(\tilde{u}_{e})=\tilde{u}_{i}$, $g(\tilde{u}_{i})=\tilde{u}_{e}$, and $F_{j}(\tilde{u}_{e})=\tilde{m}$ [^2] The stability analysis of the system without adaptation, i.e., $b=0$, is covered in Appendix A.1..
We apply the procedure from Harris & Ermentrout [31], Byrne et al [62], and Wyller et al [32]333[32], however, assumes that the fixed point values $\tilde{u}_{e}$ and $\tilde{u}_{i}$ are equal. to investigate the stability of the homogeneous steady states (i.e., the fixed points) of field models. To do so, we add a small spatial perturbation to the homogeneous steady state,
@@
Then, we insert Equations (5) into Equations (1) and linearize the system around the homogeneous steady state by applying a Taylor expansion up to the first order in $\chi(x,t)$, $\varphi(x,t)$, and $\phi(x,t)$. We obtain
@@
For $\vec{X}=\begin{pmatrix}\hat{\chi}\\ \hat{\varphi}\\ \hat{\phi}\end{pmatrix}$, Equations (6) can be written in vector notation, $\partial t\vec{X}=A_{3}(k)\vec{X}$, where
@@
$\det(A_{3}(k)-\lambda\underline{\text{I}})=$ $\lambda^{3}+\lambda^{2}(-(a_{11}(k)+a_{22}(k)+a_{33}))$
$+$ $\lambda(a_{11}(k)a_{22}(k)+a_{11}(k)a_{33}+a_{22}(k)a_{33}-a_{13}a_{31}-a_{12}(k)a_{21}(k))$
$+$ $(-a_{11}(k)a_{22}(k)a_{33}+a_{22}(k)a_{13}a_{31}+a_{33}a_{12}(k)a_{21}(k))$
$=$ $\lambda^{3}+\lambda^{2}c_{2}(k)+\lambda c_{1}(k)+c_{0}(k)$ (8)
@@
$\begin{split}c_{0}(k)&=-a_{11}(k)a_{22}(k)a_{33}+a_{22}(k)a_{13}a_{31}+a_{33}a_{12}(k)a_{21}(k)\\ c_{1}(k)&=a_{11}(k)a_{22}(k)+a_{11}(k)a_{33}+a_{22}(k)a_{33}-a_{13}a_{31}-a_{12}(k)a_{21}(k)\\ c_{2}(k)&=-(a_{11}(k)+a_{22}(k)+a_{33}).\end{split}$ (9)
Fixed points with $\alpha_{j}(k)<0$ for all $j\in\{0,1,2\}$ are stable. If the model system has a fixed point which is unstable at $k=0$ (i.e., $\alpha_{j}(0)\geq 0$ for at least one $j\in\{0,1,2\}$, see Figure 1, first row), we call the state of the system Hopf-unstable, if $\omega_{j}(0)\neq 0$, or simply unstable, if $\omega_{j}(0)=0$.
Figure 1: Examples for static (top row) and dynamic (bottom row) Turing bifurcations in a Hopf- (top row) and Turing-unstable state (bottom row) for the system with h-currents. Panels from left to right show (cf. Equations (9)): Coefficients $c_{1}(k)$ (solid line) and $c_{2}(k)$ (dashed line), coefficient $c_{0}(k)$ (solid line) and condition $c_{1}(k)c_{2}(k)-c_{0}(k)$ (dashed line), real ($\alpha_{0,1,2}(k)$, solid lines) and imaginary ($\omega_{0,1,2}(k)$, dashed lines) parts of the eigenvalues of the linearization matrix, and the corresponding activity pattern $u_{e}(x,t)$. Circles denote the wavenumbers $k_{0}$ (red) and $k_{max}$ (pink) (see text). Insets below each plot show close-ups of the thin boxes. Values for the external input currents and the adaptation parameter were $(I_{e},I_{i})=(-0.7,-0.6),\ b=-0.5$ (top row) and $(I_{e},I_{i})=(-0.3,-0.3),\ b=-0.5$ (bottom row). All other parameters are given in Table 1.
[…]
Note that there is a difference in the instability of modes between Hopf-unstable states undergoing a Turing bifurcation and other Turing-unstable states which is exemplified in Figure 1, third column (Eigenvalues). In Hopf-unstable states, unstable modes already occur for wavenumber values $0\leq k\leq k_{0}$, while in other Turing-unstable states the minimum wavenumber for unstable modes is $k_{0}$.

### v3

@@ I Introduction
One frequently used model is the Wilson-Cowan model which was first introduced in [33] in its mass model version, i.e., without a spatial domain. The mass model version displays a rich variety of emergent dynamics and has previously been used to investigate SFA-driven SOs, either with [34, 35] or without inhibition [36]. In its original form, comprising of one excitatory and one inhibitory population of neurons without SFA [37], the dynamical landscape of this model remains fairly similar when both SFA and inhibition are included, when only SFA is present, or when both are absent. A regime of oscillations (e.g., emerging due to a Hopf instability) is at the corner end of a regime of bistability, where a stable up- and down-state coexist.
@@
To include a spatial domain, Wilson and Cowan extended the mass model to a field model which is comprised of two coupled populations of excitatory and inhibitory neurons [38]. Pinto and Ermentrout [39] investigated the system without an inhibitory population, but with a linear, negative feedback that limits the networks excitability. They establish the existence of traveling pulses in such a network and connect the speed of the waves, which are induced by the slow negative feedback, to the model parameters using the Heaviside function as transfer function. A sigmoidal transfer function speeds up waves further when compared to the Heaviside transfer function. For a homogeneous, excitatory neural field model, [40] proved that lateral inhibition is required to stabilize bump solutions (i.e., stabilize the activity to not decay or vanish over time), and while [41] confirms these findings, they additionally show that local inhomogeneous input stabilizes stationary-pulse solutions in an otherwise homogeneous network. In [42], the authors investigate the same field model without inhibition in which nonlinear negative feedback to an excitatory population of neurons is produced by synaptic depression and SFA simultaneously. They identify traveling fronts and pulses, as well as stationary or homogeneous oscillations. They show that synaptic depression is not sufficient to stabilize bump solutions and confirm that either lateral inhibition (as shown in [40]) or external input (as shown in [41]) is required. Conversely, bumps do not exist when SFA is present. Furthermore, they investigated the phase space with respect to the strengths of SFA and synaptic depression and found that SFA has little to no effect on the speed and width of traveling fronts and pulses. Their nonlinear formulation of SFA feedback is a physiologically based negative feedback. In [43], the authors equip the excitatory population of the two-population model with SFA and investigate the conditions for Turing and Hopf bifurcations to emerge, enabling stationary or spatio-temporal activity. While weaker and faster adaptation generates stationary patterns, slower and stronger adaptation causes the stationary patterns to travel.
@@
Now, we are left with the question of what the differences are between the slow negative feedback mechanism based on SFA, see [44, 45, 46, 47, 35, 48], and the recently proposed positive feedback mechanism based on h-currents, see [21, 22] which have already been shown to induce SO-like dynamics. Simultaneous investigations of h-currents and SFA were already conducted. [49] used fluorescent dye in the dentate gyrus of rats and unmodified versus blocked HCN-channel [^2] HCN channels: Hyperpolarization-activated cyclic nucleotide gated channels correspond to h-current based adaptation. recordings to show that in the presence of HCN channels action potential (AP) initiation is enhanced during sustained firing and the propagation of AP patterns is sped up. They observe that in the presence of HCN channels, hyperpolarizing Na+-K+ currents are opposed. [50] included both HCN (h-current based adaptation) and M-channel (SFA) dynamics in a spiking neural network to investigate their role in seizure events in epilepsy patients. [22] found, in slice experiments, that blocking h-currents results in prolonged up- and down-state durations, effectively decreasing the temporal oscillation frequency. With their computational model they produce adaptation-driven SOs between up- and down-states where the implemented h-current and SFA based adaptation alternate in activation, depending on the membrane potential with h-currents active in down-, and inactive in up-states while the opposite is the case for SFA. With the model, they recreated the results found in vitro, prolonging up- and down-state durations, if the strength of h-currents is further reduced. [21] tested whether a spatial gradient in excitability, degree of recurrent excitation, or strength of the SFA current would induce a sequential activation of cortical neurons (i.e, up-state onset) along the dorso-ventral axis of the rat post-subiculum with a synchronous deactivation of the up-state. They observed this in vivo in naturally sleeping mice prior to their in-silico experiments. Neither gradient was able to replicate the experimentally observed dynamics. Including h-currents, on the other hand, with decreasing strength along the dorso-ventral axis replicated their results. A detailed analysis of the two opposing mechanisms shaping the spatio-temporal neural dynamics, however, has not yet been performed.
@@ II The spatially extended adaptive Wilson-Cowan model
We use the original Wilson-Cowan continuum model of [38] with integro-differential equations, including a convolution over a spatial domain, to investigate adaptation-driven traveling waves of SO-like dynamics. We include a third differential equation for an adaptation mechanism and connect it to the excitatory population to describe the effects of either high-activity induced negative or low-activity induced positive feedback mechanisms. In the following we refer to them as spike-frequency ($a$) or h-current ($h$) mediated adaptation although the model is a simplification that focuses only on their most prominent biophysical features. The equations of the activities $u_{e}$ and $u_{i}$ of the excitatory and inhibitory population and for the adaptation variable $m$ are given by
@@
The coupling strengths between the excitatory and inhibitory populations are given by connectivity weights $w_{jl},\ j,l\in\{e,i\}$, in Eq. (1). The membrane time constants are designated with $\tau_{j},\ j\in\{e,i\}$. The adaptation mechanisms $m(x,t)$ are discussed in detail in Section II.1; they are parametrized by a strength value $b$, a time constant $\tau_{m}$, and a gain parameter $\beta_{m}$. Due to the transfer function in Eq. (3) being a sigmoid between zero and one, we interpret the activity variables $u_{e}$ and $u_{i}$ as the proportion of neurons firing per unit time [54]. Its gain parameter is denoted by $\beta_{j}$. The kernel widths in Eq. (2) determine the extent to which populations are effectively connected in space. Kernel widths are chosen to be smaller for excitatory and broader for inhibitory interactions, effectively implementing a Mexican-hat shaped interaction between different locations. The parameters that are used in this study were adjusted such that adaptation-driven SOs are generated with an additional noisy external input (see [46, 34]), fast excitation-inhibition driven oscillations can emerge (see [55]), and the system with and without adaptation shows regimes of bistability since this has been shown to play a crucial part in the formation of SOs (see [51, 56]).
Table 1: List of model parameters. The abbreviations “exc" and “inh" stand for “excitatory population" and “inhibitory population", respectively.
@@
$w_{ee}$ 3.2 Coupling strength exc to exc
$w_{ei}$ 2.6 Coupling strength inh to exc
$w_{ie}$ 3.3 Coupling strength exc to inh
$w_{ii}$ 0.9 Coupling strength inh to inh
$I_{e}$ [-2.5, 2] External input current to exc
$I_{i}$ [-3.5, 1] External input current to inh
$\tau_{e}$ $10$ Membrane time constant of exc
$\tau_{i}$ $15$ Membrane time constant of inh
$\sigma_{e}$ 1 Spatial spread of exc
$\sigma_{i}$ 3 Spatial spread of inh
$\mu$ $0.4$ Adaptation threshold
@@
$b$ $[0,1]$ SFA strength
$\tau_{m}$ $300$ SFA time constant
$\beta_{m}$ $10$ SFA gain
@@
The model is implemented as a one-dimensional neural field with periodic boundary conditions and a homogeneous connectivity profile. To reduce the effects of the finite size on the dynamics, we combine a large ring of length $L=600$ with comparably small widths $\sigma_{j},\ j\in\{e,i\}$, of the spatial kernels. The ring is discretized into $n$ positions, each of which is equipped with a node whose dynamics is described by Eqs. (1). Therefore, the integration space step is given by $dx=\frac{L}{n}$.
@@ II.1 The adaptation mechanisms
The excitatory populations are equipped with an adaptation mechanism, generally understood as a process (not restricted to the brain) that modulates the effect of a sustained input but acts on a slower time scale (see [57]). In the following, we distinguish two adaptation processes caused by different types of somatic conductances in cortical neurons (see [58]). Spike-frequency adaptation (SFA), a negative feedback mechanism, is mediated by slow potassium currents which activate when the activity of the excitatory population is high. This causes a hyperpolarizing feedback current which effectively inhibits the neural response of excitatory neurons (see [47]). H-current based adaptation is mediated by hyperpolarization-activated nucleotide-gated channels (see [59]) which open if the excitatory population is in a low-activity state and return positive feedback, i.e., effectively excite the excitatory population. Functionally, the adaptation mechanisms are similar (see [60] for SFA, for h-currents see [61] where the authors discuss h-currents in thalamic neurons, and see [21, 58] for both adaptation mechanisms in cortical neurons), so that we can describe their main effects by the same differential equation (last equation in Eqs. (1), see [21]) and include the inverse mode of action (positive versus negative feedback, activated during high- versus low-activity values) by changing the sign of the parameters for adaptation strength $b$ and adaptation gain $\beta_{m}$. Parameters are summarized in Table 1.
@@ II.2 Fixed points and stability analysis of the model with adaptation
To determine the fixed points we consider a single Wilson-Cowan node. Due to the normalization of the kernel, its fixed points are equivalent to the fixed points of the spatially extended model. Let $\frac{\partial u_{e}}{\partial t}=\frac{\partial u_{i}}{\partial t}=0$ in Eqs. (1). From $\frac{\partial m}{\partial t}=0$ we obtain $m=F_{m}(u_{e})$ which we insert in the equation for the excitatory population. The nullclines $f(u_{e})$ and $g(u_{i})$ are then given by
$u_{i}=f(u_{e})=$ $\frac{1}{w_{ei}}\bigl(w_{ee}u_{e}+I_{e}-bF_{j}(u_{e})-F^{-1}_{e}(u_{e})\ \bigr),$
$j\in\{a,h\},$
@@
A point $(\tilde{u}_{e},\tilde{u}_{i},\tilde{m})\in(0,1)^{3}$ is a fixed point if $f(\tilde{u}_{e})=\tilde{u}_{i}$, $g(\tilde{u}_{i})=\tilde{u}_{e}$, and $F_{m}(\tilde{u}_{e})=\tilde{m}$ [^3] The stability analysis of the system without adaptation, i.e., $b=0$, is covered in Appendix A.1..
We apply the procedure from Harris & Ermentrout [31], Byrne et al [62], and Wyller et al [32]444[32], however, assumes that the fixed point values $\tilde{u}_{e}$ and $\tilde{u}_{i}$ are equal. to investigate the stability of the homogeneous steady states (i.e., the fixed points) of field models. To do so, we add a small spatial perturbation to the homogeneous steady state,
@@
Figure 1: Examples for static (top row) and dynamic (bottom row) Turing bifurcations in a Hopf- (top row) and Turing-unstable state (bottom row) for the system with h-currents. Panels from left to right show (cf. Eqs. (7)): Coefficients $c_{1}(k)$ (solid line) and $c_{2}(k)$ (dashed line), coefficient $c_{0}(k)$ (solid line) and condition $c_{1}(k)c_{2}(k)-c_{0}(k)$ (dashed line), real ($\alpha_{0,1,2}(k)$, solid lines) and imaginary ($\omega_{0,1,2}(k)$, dashed lines) parts of the eigenvalues of the linearization matrix, and the corresponding activity pattern $u_{e}(x,t)$. Circles denote the wavenumbers $k_{0}$ (red) and $k_{max}$ (pink) (see text). Insets below each plot show close-ups of the thin boxes. Values for the external input currents and the adaptation parameter were $(I_{e},I_{i})=(-0.7,-0.6),\ b=-0.5$ (top row) and $(I_{e},I_{i})=(-0.3,-0.3),\ b=-0.5$ (bottom row). All other parameters are given in Table 1.
Then, we insert Eqs. (4) into Eqs. (1) and linearize the system around the homogeneous steady state by applying a Taylor expansion up to the first order in $\chi(x,t)$, $\varphi(x,t)$, and $\phi(x,t)$. We obtain
@@
For $\vec{X}=\begin{pmatrix}\hat{\chi}\\ \hat{\varphi}\\ \hat{\phi}\end{pmatrix}$, Eqs. (5) can be written in vector notation, $\partial t\vec{X}=A_{3}(k)\vec{X}$, where
@@
$\det(A_{3}(k)-\lambda\underline{\text{I}})=$ $\ \lambda^{3}+\lambda^{2}\bigl(-(a_{11}(k)+a_{22}(k)+a_{33})\bigr)$
$+\lambda\bigl(a_{11}(k)a_{22}(k)+a_{11}(k)a_{33}+a_{22}(k)a_{33}-a_{13}a_{31}-a_{12}(k)a_{21}(k)\bigr)$
$+\bigl(-a_{11}(k)a_{22}(k)a_{33}+a_{22}(k)a_{13}a_{31}+a_{33}a_{12}(k)a_{21}(k)\bigr)$
$=$ $\ \lambda^{3}+\lambda^{2}c_{2}(k)+\lambda c_{1}(k)+c_{0}(k)$
@@
$\begin{split}c_{0}(k)=&\ -a_{11}(k)a_{22}(k)a_{33}+a_{22}(k)a_{13}a_{31}\\ &\ +a_{33}a_{12}(k)a_{21}(k)\\ c_{1}(k)=&\ a_{11}(k)a_{22}(k)+a_{11}(k)a_{33}+a_{22}(k)a_{33}\\ &\ -a_{13}a_{31}-a_{12}(k)a_{21}(k)\\ c_{2}(k)=&\ -(a_{11}(k)+a_{22}(k)+a_{33}).\end{split}$ (7)
Fixed points with $\alpha_{j}(k)<0$ for all $j\in\{0,1,2\}$ are stable. If the model system has a fixed point which is unstable at $k=0$ (i.e., $\alpha_{j}(0)\geq 0$ for at least one $j\in\{0,1,2\}$, see Fig. 1, first row), we call the state of the system Hopf-unstable, if $\omega_{j}(0)\neq 0$, or simply unstable, if $\omega_{j}(0)=0$.
@@
[…]
Note that there is a difference in the instability of modes between Hopf-unstable states undergoing a Turing bifurcation and other Turing-unstable states which is exemplified in Fig. 1, third column (Eigenvalues). In Hopf-unstable states, unstable modes already occur for wavenumber values $0\leq k\leq k_{0}$, while in other Turing-unstable states the minimum wavenumber for unstable modes is $k_{0}$.